The power dissipated in the 16-Ohm resistor in the
Answer #1

here,

using KVL for the loop

(16 + 1 + 1) * I = 12 - 3

I = 0.5 A

the power dissapated in 16 ohm resistor , P = I^2 * 16

P = 4 W

the power dissapated in 16 ohm resistor is 4 ohm